Solve using addition method
x+3y+z=6
3x+y-z=-2
2x+2y-z=1

Answer by Alan3354(69443)

Solve using addition method
x+3y+z=6
3x+y-z=-2
2x+2y-z=1
-----------
Add eqn 1 and eqn 2.
Then add eqn 1 and eqn 3.
---------
That eliminates z, leaves 2 equations in x & y.
Can you do the rest?